If I drive x mph for t hours and then y mph for 2t hours. Then what is expression for the average speed.

Answer by stanbon(75887)

If I drive x mph for t hours and then y mph for 2t hours. Then what is expression for the average speed.
-------------------------------
average speed = (total distance)/(total time)
-----
distance = rate*time
-------------------
Your Problem:
total distance = (xt)+ (2yt) = (x+2y)t miles
----------------------
total time = t + 2t = 3t hours
---------------------------
Therefore, average speed = [(x+2y)t]/(3t) = (x+2y)/3 mph
===============================================================
Cheers,
Stan H.